# The Leftmost particle of branching subordinators

### v1

@@ 1.3 Main results
We denote by $\kappa^{ri,\alpha,\beta}_{\cdot}(\cdot)$ the truncated Laplace exponent associated with $(0,\Lambda^{ri}_{\alpha,\beta})$. The decomposition (10)-(11) of the truncated Laplace exponent $\kappa^{ri,\alpha,\beta}_{\cdot}(\cdot)$ reads $\kappa^{ri,\alpha,\beta}_{a}(\lambda)=\phi^{\alpha}(\lambda)-M^{ri,\beta}_{a}(\lambda)$ where $\phi^{\alpha}(\lambda):=\lambda^{\alpha}$ and $M^{ri,\beta}_{a}(\lambda):=\sum_{k=2}^{\infty}e^{-\lambda(\log k)^{\beta}}\mathds{1}_{\{(\log k)^{\beta}<a\}}$. In particular, Assumption 1 is satisfied in this example. Let $\mu_{ri,\beta}$ be the measure defined by taking $\Lambda=\Lambda^{ri}_{\alpha,\beta}$ in (16). We have $\mu_{ri,\beta}([0,a))=\sharp\{k\geq 2,(\log k)^{\beta}<a\}=\lfloor e^{a^{1/\beta}}\rfloor-1$. This shows that Assumption 2 is satisfied with $\sigma:=(1-\beta)/\beta$. We denote by $M^{ri,\beta}_{2,\cdot}(\cdot)$ the function $M_{2,\cdot}(\cdot)$ associated to $\Lambda^{ri}_{\alpha,\beta}$ via (18). We have clearly $M^{ri,\beta}_{2,a}(\lambda)=(M^{ri,\beta}_{a}(\lambda))^{2}$ for any $a,\lambda\geq 0$ so, in the present example, Assumption 3 is satisfied. Since all assumptions are satisfied, we can apply Theorem 1.5 and get Corollary 1.6.
@@
For any $\ell>0$ and $a>0$ we have $M^{mb,\beta}_{a}(\ell a^{\theta-1})\leq a^{1-\theta}e^{a^{\theta}}$. For any $\ell\in(0,1/3)$ and $a>0$ we have $M^{mb,\beta}_{a}(\ell a^{\theta-1})\geq a^{1-\theta}(e^{2a^{\theta}/3}-e^{2\delta a^{\theta}/3})$ where $\delta:=(\ell+2/3)^{1/(\theta-1)}$.
  • Proof.
@@
    $M^{ri,\beta}_{a}(\lambda(a))=\int_{0}^{a}e^{\mathfrak{z}^{\theta}-\mathfrak{z}\lambda(a)}\mathrm{d}\mathfrak{z}\leq\int_{0}^{a}e^{\mathfrak{z}\lambda(a)/\ell}\mathrm{d}\mathfrak{z}=\frac{e^{\lambda(a)a/\ell}-1}{\lambda(a)/\ell}\leq a^{1-\theta}e^{a^{\theta}},$
@@
    $M^{ri,\beta}_{a}(\lambda(a))\geq\int_{\delta a}^{a}e^{2\mathfrak{z}\lambda(a)/3\ell}\mathrm{d}\mathfrak{z}=\frac{e^{2\lambda(a)a/3\ell}-e^{2\delta\lambda(a)a/3\ell}}{2\lambda(a)/3\ell}=\frac{3}{2}a^{1-\theta}(e^{2a^{\theta}/3}-e^{2\delta a^{\theta}/3}),$
    where we have used that $\mathfrak{z}^{\theta}-\mathfrak{z}\lambda(a)\geq 2\mathfrak{z}\lambda(a)/3\ell$ for all $\mathfrak{z}\in[\delta a,a]$ when $\lambda(a)=\ell a^{\theta-1}$. The second point follows. ∎
We denote by $M^{mb,\theta}_{2,\cdot}(\cdot)$ the function $M_{2,\cdot}(\cdot)$ associated to $\Lambda^{mb}_{\alpha,\theta}$ via (18). By Remark 1.4, we have $M^{mb,\theta}_{2,a}(\lambda)=M^{mb,\theta}_{a}(2\lambda)$ for any $a,\lambda\geq 0$. Therefore, taking any $c\in(0,1/3)$ and applying Lemma 2.1 (the first point with $\ell=2c$ and the second point with $\ell=c$) we get (19) so, in the present example, Assumption 3 is satisfied. Since all assumptions are satisfied, we can apply Theorem 1.5 and get Corollary 1.7.
Remark 2.2 (Trinary branchings).
In the present example, let us replace binary branchings with trinary branchings by setting that, at each branching, two particles are born (with identical displacements) instead of one. Let us denote by $\Lambda^{mt}_{\alpha,\theta}$ the corresponding jump-branching measure and by $\kappa^{mt,\alpha,\theta}_{\cdot}(\cdot)$ the truncated Laplace exponent of the branching subordinator with characteristic couple $(0,\Lambda^{mt}_{\alpha,\theta})$. The decomposition (10)-(11) of $\kappa^{mt,\alpha,\theta}_{\cdot}(\cdot)$ reads $\kappa^{mt,\alpha,\theta}_{a}(\lambda)=\phi^{\alpha}(\lambda)-M^{mt,\theta}_{a}(\lambda)$ where $\phi^{\alpha}$ is as before and $M^{mt,\theta}_{a}(\lambda)=2M^{mb,\theta}_{a}(\lambda)$. Moreover, $M^{mt,\theta}_{2,a}(\lambda)=4M^{mb,\theta}_{a}(2\lambda)$. We thus see that all the assumptions of Theorem 1.5 are identically satisfied in this case so Corollary 1.7 still holds in this case with the same value of $\gamma$.
@@ 3 Finite or infinite limit for the leftmost particle
  • Proof.
    Recall that $\mathscr{Q}^{\star}=\{\bm{x}\in\mathscr{Q},x_{2}<\infty\}$. Since $\Lambda(\mathscr{Q}^{\star})>0$ (see (3)) and $\Lambda(\{\bm{x}\in\mathscr{Q};\;x_{2}<a\})<\infty$ for any $a>0$ (see (4)), one can choose $a_{1}>0$ such that $\Lambda(\{\bm{x}\in\mathscr{Q};\;x_{2}<a_{1}\})\in(0,\infty)$. Denote by $\mathscr{Q}^{\prime}$ the set $\{\bm{x}\in\mathscr{Q};\;x_{2}<a_{1}\}$. Let $T$ be the smallest time where a branching in $\mathscr{Q}^{\prime}$ occurs for the initial particle $\varnothing$; we denote this branching by $\bm{x}(T)=(x_{n}(T))_{n\geq 1}$. Clearly, $T$ is positive and finite (by the construction from Appendix A, it follows an exponential distribution with parameter $\Lambda(\mathscr{Q}^{\prime})$) and is a stopping time. Since $\bm{x}(T)\in\mathscr{Q}^{\star}$ we have $x_{1}(T)\leq x_{2}(T)<\infty$ and, at $T$, the system contains at least two particles, one being at position $Z_{1}:=Y_{\varnothing}(T-)+x_{1}(T)$ and one at position $Z_{2}:=Y_{\varnothing}(T-)+x_{2}(T)$. We denote by $Y_{1}$ and $Y_{2}$ the systems issued from these two particles, with time shifted by $T$ and positions shifted down by $Z_{1}$ and $Z_{2}$ respectively. By the strong branching property at time $T$ (see Remark A.6), $Y_{1}$ and $Y_{2}$ are independent branching subordinators distributed as $Y$. We have $\underline{Y}(\infty)\leq(Z_{1}+\underline{Y_{1}}(\infty))\wedge(Z_{2}+\underline{Y_{2}}(\infty))$ so $\{\underline{Y}(\infty)=\infty\}\subset\{\underline{Y_{1}}(\infty)=\infty\}\cap\{\underline{Y_{2}}(\infty)=\infty\}$. We thus get $\mathbb{P}(\underline{Y}(\infty)=\infty)\leq\mathbb{P}(\underline{Y}(\infty)=\infty)^{2}$, which yields the result. ∎
@@
Since the branching subordinator $Y^{a}$ satisfies the exponential moment condition (6), we are reduced to prove the criteria under that condition, which is easier. This is done in the following lemma.
@@
  • Proof.
    The first point is a consequence of (14) (since in this case $\underline{y}>0$), itself justified in Appendix B. For any $\bm{x}\in\mathscr{Q}$ we set $N(\bm{x}):=\sharp\{n\geq 1\ \text{s.t.}\ x_{n}=0\}$ and note that (4) implies in particular that $\int_{\mathscr{Q}}(N(\bm{x})-1)_{+}\Lambda(\mathrm{d}\bm{x})<\infty$. It is easy to see from (7), monotone convergence and dominated convergence that
    $\sup_{\lambda\geq\theta}\kappa(\lambda)=\lim_{\lambda\rightarrow\infty}\kappa(\lambda)=\infty\mathds{1}_{\{d>0\}}+\Lambda(\{\bm{x}\in\mathscr{Q},N(\bm{x})=0\})-\int_{\mathscr{Q}}(N(\bm{x})-1)_{+}\Lambda(\mathrm{d}\bm{x}),$ (20)
    where $\Lambda(\{\bm{x}\in\mathscr{Q},N(\bm{x})=0\})\in[0,\infty]$. Now assume that $\sup_{\lambda\geq\theta}\kappa(\lambda)<0$. The combination of this assumption with (20) yields
    $d=0,\ \Lambda(\{\bm{x}\in\mathscr{Q},N(\bm{x})=0\})<\infty,\ \Lambda(\{\bm{x}\in\mathscr{Q},N(\bm{x})=0\})<\int_{\mathscr{Q}}(N(\bm{x})-1)_{+}\Lambda(\mathrm{d}\bm{x}).$ (21)
[…]
@@
    $\mathbb{E}[D_{1}]=\int_{\mathscr{Q}}(N(\bm{x})-1)\Lambda(dx)=\int_{\mathscr{Q}}(N(\bm{x})-1)_{+}\Lambda(dx)-\Lambda(\{\bm{x}\in\mathscr{Q},N(\bm{x})=0\})>0,$
@@
  • Proof of Proposition 1.3.
    We assume that we are in the setting of the proposition. The first claim is proved by Lemma 3.1. Let $\tilde{a}>a_{0}(\Lambda)$ be as in Lemma 3.2. We recall that the quantity $\mathscr{L}(d,\Lambda)$, defined in (12), is independent of the choice of $a>0$ so, in particular, $\mathscr{L}(d,\Lambda)=\lim_{\lambda\rightarrow\infty}\kappa_{\tilde{a}}(\lambda)$. Note that the $\tilde{a}$-truncation of $Y$, $Y^{\tilde{a}}$, satisfies the requirements of Lemma 3.3 and recall from the discussion after (11) that its Laplace exponent is $\kappa_{\tilde{a}}(\cdot)$. If $\mathscr{L}(d,\Lambda)>0$ (resp. $<0$) we have $\lim_{\lambda\rightarrow\infty}\kappa_{\tilde{a}}(\lambda)>0$ (resp. $<0$) so, by Lemma 3.3, $\mathbb{P}(\underline{Y}^{\tilde{a}}(\infty)=\infty)=1$ (resp. $=0$), and by Lemma 3.2, $\mathbb{P}(\underline{Y}(\infty)=\infty)=1$ (resp. $=0$).
@@ 3.2.1 The case $r=\rho$
Without loss of generality we assume that $r=\rho=1$. This $Y$ has characteristic couple $(0,\Lambda^{bp}_{1,1})$ and we see from Example I that $\mathscr{L}(0,\Lambda^{bp}_{1,1})=0$. Defining $(C_{t})_{t\geq 0}$ and $(D_{t})_{t\geq 0}$ as in the proof of Lemma 3.3 we have that (22) still holds true but this time $\mathbb{E}[D_{1}]=\int_{\mathscr{Q}}(N(\bm{x})-1)\Lambda(dx)=0$. Since $(D_{t})_{t\geq 0}$ is almost surely not constant we get that $(D_{t})_{t\geq 0}$ oscillates, and since $(D_{t})_{t\geq 0}$ only take integer values and only has jumps of size $1$ and $-1$, we get that $\mathbb{P}(\forall t\geq 0,D_{t}\neq 0)=0$. Therefore, by (22) we get $\mathbb{P}(\underline{Y}(\infty)=0)=0$. Now let us prove by induction on $k$ that $\mathbb{P}(\underline{Y}(\infty)=k)=0$ for all $k\geq 0$. The claim has been proved for $k=0$ so we now fix $k\geq 0$ and assume that the claim has been proved for all $j\in\{0,\cdots,k\}$. Let $T:=\inf\{t\geq\ \text{s.t.}\ \underline{Y}(t)=k+1\}$. This $T$ is a stopping time and, by the induction hypothesis, we have almost surely $T<\infty$. Let $M:=\sharp\{j\geq 1,Y_{j}(T)=k+1\}$ and let us denote by $Y_{1},\cdots,Y_{M}$ the systems issued from the $M$ particles that are located at $k+1$ at time $T$, shifted down by $k+1$. By the strong branching property at time $T$ (see Remark A.6), we have that, for any $m\geq 1$, conditionally on $\{M=m\}$, $Y_{1},\cdots,Y_{m}$ are $m$ independent branching subordinators distributed as $Y$. In particular,
@@
We thus have almost sure convergence of $\underline{Y}(t)$ to infinity as $t$ goes to infinity, but we see from (14) that $\underline{Y}(t)/t$ converges to 00. The branching Poisson process in the case $r=\rho$ thus provides another instance of sub-linear growth of $\underline{Y}(t)$, even if it does not satisfy the assumptions from Theorem 1.5.

### v2

@@ 1.3.4 Organization of the paper
We denote by $\kappa^{ri,\alpha,\beta}_{\cdot}(\cdot)$ the truncated Laplace exponent associated with $(0,\Lambda^{ri}_{\alpha,\beta})$. The decomposition (11)-(12) of the truncated Laplace exponent $\kappa^{ri,\alpha,\beta}_{\cdot}(\cdot)$ reads $\kappa^{ri,\alpha,\beta}_{a}(\lambda)=\phi^{\alpha}(\lambda)-M^{ri,\beta}_{a}(\lambda)$ where $\phi^{\alpha}(\lambda):=\lambda^{\alpha}$ and $M^{ri,\beta}_{a}(\lambda):=\sum_{k=2}^{\infty}e^{-\lambda(\log k)^{\beta}}\mathds{1}_{\{(\log k)^{\beta}<a\}}$. In particular, Assumption 1 is satisfied in this example. Let $\mu_{ri,\beta}$ be the measure defined by taking $\Lambda=\Lambda^{ri}_{\alpha,\beta}$ in (20). We have $\mu_{ri,\beta}([0,a))=\sharp\{k\geq 2;(\log k)^{\beta}<a\}=\lfloor e^{a^{1/\beta}}\rfloor-1$. This shows that Assumption 2 is satisfied with $\sigma:=(1-\beta)/\beta$. We denote by $M^{ri,\beta}_{2,\cdot}(\cdot)$ the function $M_{2,\cdot}(\cdot)$ associated to $\Lambda^{ri}_{\alpha,\beta}$ via (22). We have clearly $M^{ri,\beta}_{2,a}(\lambda)=(M^{ri,\beta}_{a}(\lambda))^{2}$ for any $a,\lambda\geq 0$ so, in the present example, Assumption 3 is satisfied. Finally, $\Lambda^{ri}_{\alpha,\beta}$ clearly satisfies Assumption 4. Since all assumptions are satisfied, we can apply Theorem 1.5 and get Corollary 1.6.
@@
For any $\ell>0$ and $a>0$ we have $M^{mb,\theta}_{a}(\ell a^{\theta-1})\leq a^{1-\theta}e^{a^{\theta}}$. For any $\ell\in(0,1/3)$ and $a>0$ we have $M^{mb,\theta}_{a}(\ell a^{\theta-1})\geq a^{1-\theta}(e^{2a^{\theta}/3}-e^{2\delta a^{\theta}/3})$ where $\delta:=(\ell+2/3)^{1/(\theta-1)}$.
@@
$M^{mb,\theta}_{a}(\lambda(a))=\int_{0}^{a}e^{\mathfrak{z}^{\theta}-\mathfrak{z}\lambda(a)}\mathrm{d}\mathfrak{z}\leq\int_{0}^{a}e^{\mathfrak{z}\lambda(a)/\ell}\mathrm{d}\mathfrak{z}=\frac{e^{\lambda(a)a/\ell}-1}{\lambda(a)/\ell}\leq a^{1-\theta}e^{a^{\theta}},$
@@
$M^{mb,\theta}_{a}(\lambda(a))\geq\int_{\delta a}^{a}e^{2\mathfrak{z}\lambda(a)/3\ell}\mathrm{d}\mathfrak{z}=\frac{e^{2\lambda(a)a/3\ell}-e^{2\delta\lambda(a)a/3\ell}}{2\lambda(a)/3\ell}=\frac{3}{2}a^{1-\theta}(e^{2a^{\theta}/3}-e^{2\delta a^{\theta}/3}),$
where we have used that $\mathfrak{z}^{\theta}-\mathfrak{z}\lambda(a)\geq 2\mathfrak{z}\lambda(a)/3\ell$ for all $\mathfrak{z}\in[\delta a,a]$ when $\lambda(a)=\ell a^{\theta-1}$. The second point follows. ∎ We denote by $M^{mb,\theta}_{2,\cdot}(\cdot)$ the function $M_{2,\cdot}(\cdot)$ associated to $\Lambda^{mb}_{\alpha,\theta}$ via (22). By Remark 1.4, we have $M^{mb,\theta}_{2,a}(\lambda)=M^{mb,\theta}_{a}(2\lambda)$ for any $a,\lambda\geq 0$. Therefore, taking any $c\in(0,1/3)$ and applying Lemma 2.1 (the first point with $\ell=2c$ and the second point with $\ell=c$) we get (23) so, in the present example, Assumption 3 is satisfied. Finally, $\Lambda^{mb}_{\alpha,\theta}$ clearly satisfies Assumption 4. Since all assumptions are satisfied, we can apply Theorem 1.5 and get Corollary 1.7.
Remark 2.2 ($p$-ary branchings).
In the present example, let us set $p\geq 3$ and replace binary branchings with $p$-ary branchings by setting that, at each branching, p1p-1 particles are born (with identical displacements) instead of one. Let us denote by $\Lambda^{mp}_{\alpha,\theta}$ the corresponding jump-branching measure and by $\kappa^{mp,\alpha,\theta}_{\cdot}(\cdot)$ the truncated Laplace exponent of the branching subordinator with characteristic couple $(0,\Lambda^{mp}_{\alpha,\theta})$. The decomposition (11)-(12) of $\kappa^{mp,\alpha,\theta}_{\cdot}(\cdot)$ reads $\kappa^{mp,\alpha,\theta}_{a}(\lambda)=\phi^{\alpha}(\lambda)-M^{mp,\theta}_{a}(\lambda)$ where $\phi^{\alpha}$ is as before and $M^{mp,\theta}_{a}(\lambda)=(p-1)M^{mb,\theta}_{a}(\lambda)$. Moreover, $M^{mp,\theta}_{2,a}(\lambda)=(p-1)^{2}M^{mb,\theta}_{a}(2\lambda)$. We thus see that all the assumptions of Theorem 1.5 are identically satisfied in this case so Corollary 1.7 still holds in this case with the same value of $\gamma$.
@@ 3 Finite or infinite limit for the leftmost particle
To prove Lemma 3.1 we use that the property of the leftmost particle converging to $\infty$ is inherited in the sense that the property is also satisfied by all subsystems formed by one particle and its future lineage. It is a classical fact that, for Galton-Watson trees, any inherited property occurs with probability 0 or $1$, conditionally on non-extinction, see Proposition 5.6 in [23]. The following proof adapts this simple idea to branching subordinators.
Recall that $\mathscr{Q}^{\star}=\{\bm{x}\in\mathscr{Q};x_{2}<\infty\}$. Since $\Lambda(\mathscr{Q}^{\star})>0$ (see (3)) and $\Lambda(\{\bm{x}\in\mathscr{Q};\;x_{2}<a\})<\infty$ for any $a>0$ (see (4)), one can choose $a_{1}>0$ such that $\Lambda(\{\bm{x}\in\mathscr{Q};\;x_{2}<a_{1}\})\in(0,\infty)$. Denote by $\mathscr{Q}^{\prime}$ the set $\{\bm{x}\in\mathscr{Q};\;x_{2}<a_{1}\}$. Let $T$ be the smallest time where a branching in $\mathscr{Q}^{\prime}$ occurs for the initial particle $\varnothing$; we denote this branching by $\bm{x}(T)=(x_{n}(T))_{n\geq 1}$. Clearly, $T$ is positive and finite (by the construction from Appendix A, it follows an exponential distribution with parameter $\Lambda(\mathscr{Q}^{\prime})$) and is a stopping time. Since $\bm{x}(T)\in\mathscr{Q}^{\star}$ we have $x_{1}(T)\leq x_{2}(T)<\infty$ and, at $T$, the system contains at least two particles, one being at position $Z_{1}:=Y_{\varnothing}(T-)+x_{1}(T)$ and one at position $Z_{2}:=Y_{\varnothing}(T-)+x_{2}(T)$. We denote by $Y_{1}$ and $Y_{2}$ the systems issued from these two particles, with time shifted by $T$ and positions shifted down by $Z_{1}$ and $Z_{2}$ respectively. By the strong branching property at time $T$ (see Remark A.6), $Y_{1}$ and $Y_{2}$ are independent branching subordinators distributed as $Y$. We have $\underline{Y}(\infty)\leq(Z_{1}+\underline{Y_{1}}(\infty))\wedge(Z_{2}+\underline{Y_{2}}(\infty))$ so $\{\underline{Y}(\infty)=\infty\}\subset\{\underline{Y_{1}}(\infty)=\infty\}\cap\{\underline{Y_{2}}(\infty)=\infty\}$. We thus get $\mathbb{P}(\underline{Y}(\infty)=\infty)\leq\mathbb{P}(\underline{Y}(\infty)=\infty)^{2}$, which yields the result. ∎
@@
Since the branching subordinator $Y^{a}$ satisfies the exponential moment condition (6) (see Section 1.2), we are reduced to prove the criterion under that condition. This is done in the following lemma.
@@
First assume that $\sup_{\lambda\geq\theta}\kappa(\lambda)>0$. We can thus fix $\lambda\geq\theta$ such that $\kappa(\lambda)>0$. Then, using (8),
$e^{-\lambda\underline{Y}(t)}\leq\sum_{u\in\mathscr{N}(t)}e^{-\lambda Y_{u}(t)}=e^{-t\kappa(\lambda)}W_{\lambda}(t).$
Combining with the almost sure convergence of the non-negative martingale $(W_{\lambda}(t))_{t\geq 0}$ we get that, almost surely, $e^{-\lambda\underline{Y}(t)}$ converges to 0 as $t\to\infty$, so $\mathbb{P}(\underline{Y}(\infty)=\infty)=1$.
Now assume that $\sup_{\lambda\geq\theta}\kappa(\lambda)<0$. Combining this assumption with (16) and recalling, from the discussion before Remark 1.2, that $\int_{\mathscr{Q}}(N(\bm{x})-1)_{+}\Lambda(\mathrm{d}\bm{x})<\infty$, we get
$d=0,\ \Lambda(\{\bm{x}\in\mathscr{Q};N(\bm{x})=0\})<\int_{\mathscr{Q}}(N(\bm{x})-1)_{+}\Lambda(\mathrm{d}\bm{x})<\infty.$ (24)
Let $C_{t}:=\sum_{u\in\mathscr{N}(t)}\mathds{1}_{\{Y_{u}(t)=0\}}$. At time $t$, the jump rate of the process $C$ is $C_{t}\Lambda(\{\bm{x}\in\mathscr{Q};N(\bm{x})\neq 1\})$ (which is finite by (24)) and the size of jumps has the law of $N(\bm{x})-1$ under $\Lambda(\{\bm{x}\in\mathscr{Q};N(\bm{x})\neq 1\}\cap\cdot)/\Lambda(\{\bm{x}\in\mathscr{Q};N(\bm{x})\neq 1\})$. We have $C_{0}=1$ and $(C_{t})_{t\geq 0}$ remains constant equal to 0 after reaching 0 for the first time. Let $(D_{t})_{t\geq 0}$ be a compound Poisson process starting from $1$, that jumps at rate $\Lambda(\{\bm{x}\in\mathscr{Q};N(\bm{x})\neq 1\})$ and whose jumps have the same law as those of $(C_{t})_{t\geq 0}$. Let $I_{t}:=\int_{0}^{t}(1/D_{s})ds$ with the convention $1/0=\infty$. Let $A_{t}:=\inf\{s\geq 0;\ I_{s}>t\}$. A classical argument shows that $(C_{t})_{t\geq 0}$ is equal in law to $(D_{A_{t}})_{t\geq 0}$. Therefore,
@@
$\mathbb{E}[D_{1}]=\int_{\mathscr{Q}}(N(\bm{x})-1)\Lambda(\mathrm{d}\bm{x})=\int_{\mathscr{Q}}(N(\bm{x})-1)_{+}\Lambda(\mathrm{d}\bm{x})-\Lambda(\{\bm{x}\in\mathscr{Q};N(\bm{x})=0\})>0,$
@@
Remark 3.4.
The proof of the first point of Lemma 3.3 relies on the existence of the Laplace exponent $\kappa(\cdot)$ and therefore requires Assumption (6). However, using (15) instead of (16), the proof of the second point applies under the less restrictive assumptions of Proposition 1.3 and allows to yields directly the second point of that proposition.
Remark 3.5.
Under the extra assumption that $\infty<\kappa(\theta_{0})<0$ for some $\theta_{0}>0$, the first point of Lemma 3.3 is also a consequence of (18) since in this case $\underline{y}>0$.
@@
[…]
@@ 3.2.1 The case $r=\rho$
Without loss of generality we assume that $r=\rho=1$. This $Y$ has characteristic couple $(0,\Lambda^{bp}_{1,1})$ and we see from Example I that $\mathscr{L}(0,\Lambda^{bp}_{1,1})=0$. Defining $(C_{t})_{t\geq 0}$ and $(D_{t})_{t\geq 0}$ as in the proof of Lemma 3.3 we have that (25) still holds true but this time $\mathbb{E}[D_{1}]=\int_{\mathscr{Q}}(N(\bm{x})-1)\Lambda(\mathrm{d}\bm{x})=0$. Since the Lévy process $(D_{t})_{t\geq 0}$ is almost surely non-constant we get that $(D_{t})_{t\geq 0}$ oscillates, and since that process only takes integer values and only has jumps of size $1$ and 1-1, we get that $\mathbb{P}(\forall t\geq 0,D_{t}\neq 0)=0$. Therefore, by (25) we get $\mathbb{P}(\underline{Y}(\infty)=0)=0$. Now let us prove by induction on $k$ that $\mathbb{P}(\underline{Y}(\infty)=k)=0$ for all $k\geq 0$. The claim has been proved for $k=0$ so we now fix $k\geq 0$ and assume that the claim has been proved for all $j\in\{0,\cdots,k\}$. Let $T:=\inf\{t\geq 0;\ \underline{Y}(t)=k+1\}$. This $T$ is a stopping time and, by the induction hypothesis, we have almost surely $T<\infty$. Let $M:=\sharp\{u\in\mathscr{N}(T);Y_{u}(T)=k+1\}$ and let us denote by $Y_{1},\cdots,Y_{M}$ the systems issued from the $M$ particles that are located at $k+1$ at time $T$, shifted down by $k+1$. By the strong branching property at time $T$ (see Remark A.6), we have that, for any $m\geq 1$, conditionally on $\{M=m\}$, $Y_{1},\cdots,Y_{m}$ are $m$ independent branching subordinators distributed as $Y$. In particular,
@@
We thus have almost sure convergence of $\underline{Y}(t)$ to infinity as $t$ goes to infinity, but we see from (18) that $\underline{Y}(t)/t$ converges to 0. The branching Poisson process in the case $r=\rho$ thus provides another instance of sub-linear growth of $\underline{Y}(t)$, even if it does not satisfy the assumptions from Theorem 1.5.